### v1

@@ 1 Introduction
Machine-learning models are used in many different contexts within an experiment [3, 4, 5, 6, 7]. Some of these uses, such as $b$-tagging or electron/photon identification, are attempting to reconstruct a well-defined property of the physical event and can be relatively well-modelled using efficiency functions. In other cases, however, the ML model defines the analysis output, for example a signal-vs-background score, where the reusability of the analysis directly depends on the reinterpretability of the ML model. There also exist cases in between (e.g. an analysis-specific top-tagger trained specifically on signal samples, where the efficiencies may differ significantly for different signal models), and all these use cases can take a wide variety of different inputs requiring different amounts of detector simulation/emulation.
@@ 2 Mechanisms and examples
Standards exist for a degree of ML interoperability. For neural networks (NNs), the most common format is ONNX [11], though, within ATLAS in particular, the LWTNN [12] format is also used. For boosted decision trees (BDTs), possible output formats include dependency-free serialisation through petrify-bdt [13] and treelite [14], custom formats like XML files from ROOT TMVA [15, 16], and to some extent also ONNX. Which output formats are available is likely to depend on the training environment, meaning that consideration to the preservation strategy is needed from an early stage of the ML-analysis design. There also exist ML architectures for which there is no current framework-independent representation, such as elements of generative networks used in normalizing-flow methods: this challenge is hence one that can be expected to continue, and the practicalities of preservation and reuse need to be considered when choosing an ML architecture.
@@
  • The ATLAS search for $R$-parity-violating supersymmetry, ATLAS-SUSY-2019-04 [22] published trained networks in ONNX format on HEPData [23]. This analysis uses 64 high-level variables plus a continuous $b$-tag score from an ML-based tagger. Also in this case, a truth-level implementation is available [24], which proved essential for understanding the input variables. However, it is unclear how to reproduce the $b$-tag score, and how to verify that physics objects from any given fast-simulation package produce the intended ML responses within acceptable uncertainties.
@@ 3 Analysis design
We note, however, that newer techniques such as normalizing flows (NFs) are gaining significant uptake for publishing likelihoods [50] and many other applications [51]. However, currently, NFs are not supported by ONNX, and active involvement from the HEP community might be necessary to achieve this. This reflects the constantly changing nature of this technology and its format ecosystem. This will likely evolve further, adding a requirement that HEP archival systems perform periodic re-exporting, versioning, and regression testing of analysis ONNX data files to maintain usability with the evolving tool set. (As with many issues in ML reinterpretation, this concern also applies to other archive formats such as HistFactory, HS3, or ROOT, but the relative complexity and novelty of ML methods make the issue particularly acute.)
@@ 5 Surrogate models
  • Particle-level surrogates: 4-vectors of all particle-level objects after showering and hadronisation (with or without detector smearing), i.e. hadrons, leptons, photons or long-lived BSM particles. This level may also include information about the positions and timings of decay or production vertices.
@@ Acknowledgements
AB acknowledges support by the UK STFC Consolidated Grant scheme (ST/W000520/1) and TP an STFC doctoral training studentship. The work of SK is supported in part by the IN2P3 master project “Théorie – DATAMatter”; financial support by the CNRS Formation Permanente for attending the Les Houches workshop is also gratefully acknowledged. AK and AR are supported by the Research Council of Norway (RCN) through the FRIPRO grant 323985 PLUMBIN’. AL is supported by FAPESP grants no. 2018/25225-9 and 2021/01089-1. HRG is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257: Particle Physics Phenomenology after the Higgs Discovery. HRG also acknowledges the support from the Italian PRIN grant 20172LNEEZ. KR is supported by the Norwegian Financial Mechanism 2014-2021, grant no. 2019/34/H/ST2/00707; and by the National Science Centre, Poland grant 2019/35/B/ST2/02008. The work of SS is supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education under contract NRF-2021R1I1A3048138.

### v3

@@ 1 Introduction
Machine-learning models are used in many different contexts within an experiment [3, 4, 5, 6, 7]. Some of these uses, such as $b$-tagging or electron/photon identification, are attempting to reconstruct a well-defined property of the physical event and can be relatively well-modelled using efficiency functions. In other cases, however, the ML model defines the analysis output, for example a signal-vs-background score, so that reinterpretability of the analysis directly depends on the reusability of the ML model. [^1] Throughout this document by reinterpretability we mean reusability of the ML model or the analysis to infer results for scenarios not originally considered by the experimental analysis. There also exist cases in between (e.g. an analysis-specific top-tagger trained specifically on signal samples, where the efficiencies may differ significantly for different signal models), and all these use cases can take a wide variety of different inputs requiring different amounts of detector simulation/emulation.
@@ 2 Mechanisms and examples
Standards exist for a degree of ML interoperability. For neural networks (NNs), the most common format is Open Neural Network Exchange (ONNX) [11], though, within ATLAS in particular, the Light-Weight Neural Network (LWTNN) [12] format is also used. For boosted decision trees (BDTs), possible output formats include dependency-free serialisation through petrify-bdt [13] and treelite [14], custom formats like XML files from ROOT TMVA [^2] Toolkit for Multivariate Data Analysis with ROOT. [15, 16], and to some extent also ONNX. Which output formats are available is likely to depend on the training environment, meaning that consideration to the preservation strategy is needed from an early stage of the ML-analysis design. There also exist ML architectures for which there is no current framework-independent representation, such as elements of generative networks used in normalizing-flow methods: this challenge is hence one that can be expected to continue, and the practicalities of preservation and reuse need to be considered when choosing an ML architecture.
@@
  • The ATLAS search for $R$-parity-violating supersymmetry, ATLAS-SUSY-2019-04 [22] published trained networks in ONNX format on HEPData [23]. This analysis uses 64 high-level variables plus a continuous $b$-tag score from a ML-based tagger. Also in this case, a truth-level implementation is available [24], which proved essential for understanding the input variables. However, it is unclear how to reproduce the $b$-tag score, and how to verify that physics objects from any given fast-simulation package produce the intended ML responses within acceptable uncertainties.
@@ 3 Analysis design
We note, however, that newer techniques such as normalizing flows (NFs) are gaining significant uptake for publishing likelihoods [50] and many other applications [51]. However, currently, NFs are not supported by ONNX, and active involvement from the HEP community might be necessary to achieve this. This reflects the constantly changing nature of this technology and its format ecosystem. This will likely evolve further, adding a requirement that HEP archival systems perform periodic re-exporting, versioning, and regression testing of analysis ONNX data files to maintain usability with the evolving tool set. (As with many issues in ML reinterpretation, this concern also applies to other archive formats such as HistFactory [52], HS3  [^3] High Energy Physics Statistics Serialization Standard. [53], or ROOT, but the relative complexity and novelty of ML methods make the issue particularly acute.)
@@ 5 Surrogate models
  • Particle-level surrogates: 4-vectors of all particle-level objects after showering and hadronisation (with or without detector smearing), i.e. hadrons, leptons, photons or long-lived Beyond-the-Standard Model (BSM) particles. This level may also include information about the positions and timings of decay or production vertices.
@@ Acknowledgements
AB acknowledges support by the UK STFC Consolidated Grant scheme (ST/W000520/1) and TP an STFC doctoral training studentship. The work of JK is supported by the Alexander von Humboldt-Stiftung. The work of SK is supported in part by the IN2P3 master project “Théorie – DATAMatter”; financial support by the CNRS Formation Permanente for attending the Les Houches workshop is also gratefully acknowledged. AK and AR are supported by the Research Council of Norway (RCN) through the FRIPRO grant 323985 PLUMBIN’. AL is supported by FAPESP grants no. 2018/25225-9 and 2021/01089-1. HRG is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant 396021762 – TRR 257: Particle Physics Phenomenology after the Higgs Discovery. HRG also acknowledges the support from the Italian PRIN grant 20172LNEEZ. KR is supported by the Norwegian Financial Mechanism 2014-2021, grant no. 2019/34/H/ST2/00707; and by the National Science Centre, Poland grant 2019/35/B/ST2/02008. The work of SS is supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education under contract NRF-2021R1I1A3048138.